f(x)=x^2-7 I'm supposed tell weather the function has a minimum or a maximum value, and we're supposed to find the minumum and maxmum value as well. I donot know how should i procced with the problem. Do I solve the problem with the formula
x=-b/2a yes or no?
Thank You very much for taking time out of your day to help me with my Algebra 1 homework, I really apperciate it. Have a nice day, and thank you again.

Answer by dabanfield(803)

f(x) = x^2 -7 is a parabola.
the max or min of a parabola is always the vertex
rewrite the equation in "parabolic vertex for
That is:
f(x) = a(x-h)²+k
where (h,k) is the vertex
and a determines how wide or narrow the parabola is, and whether it opens up or down
if a is negative it opens down, and you have a max
if a is positive then it opens up and you have a min
In this case the equation in standard form is:
f(x) = 1*(x-0)^2 + (-7)
The vertex then is (0,-7). Since a = 1 is positive the vertex is at a minimum point.